OpenStudy (anonymous):

A $200.00 sweater has been marked down to $120.00. What percentage discount from the original price does the markdown represent?

OpenStudy (anonymous):

200*x=120 where x is the percentage mark down. Solve for x

OpenStudy (anonymous):

60%

OpenStudy (ash2326):

@brittanynhall before we find the percentage discount. let's find the discount in $ discount= $200-$120 what do you get?

OpenStudy (anonymous):

180

OpenStudy (ash2326):

\( 200-120\ne 180\)

OpenStudy (ash2326):

so we get discount as $80 discount in %= \(\large \frac{80}{200}\times 100=??\)

OpenStudy (anonymous):

Do poportion

OpenStudy (anonymous):

Cost /total =x/100 Then cross multiply